# A dispersive approach to the CP conserving $K\to\pi\ell^{+}\ell^{-}$ radiative decays

### v1

@@ 1 Introduction
Concerning the charged amplitude $K^{\pm}\to\pi^{\pm}\ell^{+}\ell^{-}$, several experiments have performed measurements of both the branching fraction and the decay distribution as a function of the dilepton energy $s$, in the $e^{+}e^{-}$ [3, 4] as well as in the $\mu^{+}\mu^{-}$ mode [5, 6]. The most recent results by the NA62 collaboration are based on a sample of $28000$ events in the muon mode. The collaboration expects to collect over 100K events in each of the $\ell^{+}\ell^{-}$ modes at the end of the present run.
@@
Arguments have been proposed which favour a positive sign for the interference term between the direct and indirect CP-violating contributions to the $K_{L}\to{\pi^{0}}\ell^{+}\ell^{-}$ amplitude [11, 12]. More recently, it has been shown [13] that a positive interference is actually a property of QCD in the limit where the number of colours $N_{c}$ becomes infinite. In principle, lattice QCD is the most appropriate approach for evaluating this sign from first principles, but the present results on the radiative amplitudes are not yet sufficiently precise [14]. In this context, it would be useful to determine both the magnitude and the (appropriately defined [^1] We will define this phase here relative to the (non-perturbative) chiral coupling $G_{8}$. In order to correctly compute the interference term in the $K_{L}\to{\pi^{0}}\ell^{+}\ell^{-}$ branching fraction knowledge of the relative phase between $G_{8}$ and the imaginary part of perturbative parameter $C_{7V}$ is also necessary.) phase of the $K_{S}\to{\pi^{0}}\ell^{+}\ell^{-}$ amplitude from experiment, given a theoretical framework for analysing the data. The framework that has been usually employed is the so-called “beyond one-loop model” (B1L) proposed in ref. [15]. As compared with the calculation in the chiral expansion at order $p^{4}$ [16, 17] the B1L model includes terms of higher chiral order, which are needed in order to properly describe the experimental energy dependence. The B1L amplitudes are written as follows
@@ 3.1 Asymptotic behaviour of $W^{[1/2]}(s;\nu)$ and $W^{[3/2]}(s;\nu)$
$\lim_{q\to\infty}\int d^{4}x\,e^{iqx}T\{j_{\mu}^{em}(x)\mathcal{L}_{\Delta{S}=1}(0)\}=(q_{\mu}q_{\nu}-g_{\mu\nu}q^{2})(A+B\ln\frac{-q^{2}}{\nu^{2}}+\cdots)\,[\bar{s}\gamma_{\nu}(1-\gamma^{5}){d}].$ (17)
@@
${\displaystyle\int}d^{4}x\,e^{iq\cdot x}T\{j^{\mu}(x){\cal L}^{eff}_{\Delta{S=1}}(0)\}_{\rm dim=6}\sim\left[C^{(6)}_{V}(\nu)\dfrac{g^{\mu\beta}q^{\alpha}-g^{\mu\alpha}q^{\beta}}{q^{2}}+C^{(6)}_{A}(\nu)\epsilon^{\mu\nu\alpha\beta}\dfrac{q_{\nu}}{q^{2}}\right](\bar{s}\gamma_{\alpha}{u})(\bar{u}\gamma_{\beta}{d}).$ (18)
@@ 3.2 Muskhelishvili-Omnès representations
$\Omega^{I}_{1}(s)=\exp\left[\frac{s}{\pi}\int_{(m_{\pi}+m_{K})^{2}}^{\infty}ds^{\prime}\dfrac{\delta_{1}^{I}(s^{\prime})}{s^{\prime}(s^{\prime}-s)}\right]\$(19)
The $K\pi$ phases $\delta_{1}^{I}(s)$ to be used in these integrals must be equal to the elastic scattering phase-shifts in the elastic region and can be chosen in a somewhat arbitrary way above that region. It is noteworthy that the elastic region extends effectively well above the first inelastic threshold $K\pi\pi$, up to $\Lambda=1.3$ GeV approximately [35]. The $J=1$, $I=3/2$ phase-shift is very small and will be simply interpolated to zero above 1 GeV. In the case of $I=1/2$ we can use the $K\pi$ vector form factor (the modulus of which is a measurable quantity) i.e. take $\Omega_{1}^{1/2}(s)\equiv f_{+}^{K\pi}(s)/f_{+}^{K\pi}(0)$ which obeys eq. (19) with a phase given by the $K\pi$ form factor phase in the inelastic region.
@@
$W^{[1/2]}(s;\nu)=W^{[1/2]}_{\pi\pi}(s)+f_{+}^{K\pi}(s)\Big[C^{[1/2]}(\nu)-\frac{s}{\pi}\int_{(m_{K}+m_{\pi})^{2}}^{\Lambda^{2}}ds^{\prime}\frac{W^{[1/2]}_{\pi\pi}(s^{\prime}){\rm Im\,}[1/f_{+}^{K\pi}(s^{\prime})]}{s^{\prime}(s^{\prime}-s)}\Big]\ .$ (24)
@@
$W^{[1/2]}(s)=W^{[1/2]}_{\pi\pi}(s)+f_{+}^{K\pi}(s)\Big[G_{F}m_{K}^{2}\frac{2a_{+}-a_{S}}{f_{+}^{K\pi}(0)}-\frac{s}{\pi}\int_{(m_{K}+m_{\pi})^{2}}^{\Lambda^{2}}ds^{\prime}\frac{W^{[1/2]}_{\pi\pi}(s^{\prime}){\rm Im\,}[1/f_{+}^{K\pi}(s^{\prime})]}{s^{\prime}(s^{\prime}-s)}\Big]\ .$ (25)
@@ 4.2 Input for the $K\pi\to\pi\pi$ amplitudes
$\widetilde{M}_{1}(s)=\tilde{\mu}_{1}\widetilde{{\cal S}}_{1}(s)\ ,\quad\widehat{\widetilde{M}}_{1}(s)=\frac{\tilde{\mu}_{1}}{2\kappa(s)}\int_{-1}^{+1}dz(-9(s-s_{0})z-3\kappa(s)z^{2})\widetilde{S}_{1}(t(z))\ .$ (35)
@@ 5.5 Additional contributions: $\omega$ and $\phi$ resonances
$\bar{b}_{+}|_{\omega,\phi}\simeq+0.25\ (a_{S}>0),\quad\bar{b}_{+}|_{\omega,\phi}\simeq+0.26\ (a_{S}<0)\ .$ (54)
This reflects the fact that ${\rm Im\,}[W_{+}]$ (unlike ${\rm Re\,}[W_{+}]$), from which the combination $2g_{V}^{5}-g_{V}^{6}$ is derived, has a very small dependence on $a_{S}$ at low energy and up to the $\rho$ mass.
@@
The results of including the effects of the $I=0$ resonances, estimated in the way described above, is illustrated in fig. 13. In the case of $W_{+}$, their effect is to decrease slightly the modulus squared of the amplitudes in both cases of $a_{S}>0$ and $a_{S}<0$. The former case thus remains favoured by comparison with the experimental data. A suppressed contribution from the $I=0$ resonances to $W_{+}$ was observed also in the factorisation approach [27, 63]. In the case of $|W_{S}|^{2}$, the effect of the $I=0$ resonances is also rather small but it tends to reduce the difference between the results corresponding to $a_{S}>0$ and $a_{S}<0$ (see fig. 13, right).
@@ 5.6 Additional contributions: higher mass resonances
$C^{[1/2]}(s;\nu)=\frac{G_{F}}{\sqrt{2}}V_{ud}V_{us}\frac{8m_{K}^{2}}{3}\,(N_{c}(C_{1}-C_{4})+C_{2}-C_{3})\Big(\frac{5}{3}+\log\frac{\nu^{2}}{M^{2}}-\psi(3-\frac{s}{M^{2}})\,\Big)$ (55)
@@ 6 Conclusions
Other light vector resonances are present, in particular $\omega(782),\phi(1019)$, but their influence on the $W_{+,S}$ amplitudes cannot be accounted for in the same way as was done for the $\rho$ and the $K^{*}$ resonances, because the corresponding discontinuities involve amplitudes, $\pi\to 3\pi, KK$${\bar{K}}\to{K}\pi$ which cannot be accessed experimentally. We estimated their contributions in an approximate way by appealing to nonet symmetry, $\Delta{I}=1/2$ dominance, and relying on a chiral-resonance Lagrangian, finding them to be much less important than those of the $I=1$ and $I=1/2$ resonances.

### v2

@@ 1 Introduction
Concerning the charged amplitude $K^{\pm}\to\pi^{\pm}\ell^{+}\ell^{-}$, several experiments have performed measurements of both the branching fraction and the decay distribution as a function of the dilepton energy $s$, in the $e^{+}e^{-}$ [3, 4] as well as in the $\mu^{+}\mu^{-}$ mode [5, 6]. Comparing the $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ amplitudes was shown to provide constraints on a class of operators which violate lepton flavour universality [7, 8]. The most recent results by the NA62 collaboration are based on a sample of $28000$ events in the muon mode. The collaboration expects to collect over 100K events in each of the $\ell^{+}\ell^{-}$ modes at the end of the present run.
@@
Arguments have been proposed which favour a positive sign for the interference term between the direct and indirect CP-violating contributions to the $K_{L}\to{\pi^{0}}\ell^{+}\ell^{-}$ amplitude [13, 14]. More recently, it has been shown [15] that a positive interference is actually a property of QCD in the limit where the number of colours $N_{c}$ becomes infinite. In principle, lattice QCD is the most appropriate approach for evaluating this sign from first principles, but the present results on the radiative amplitudes are not yet sufficiently precise [16]. In this context, it would be useful to determine both the magnitude and the (appropriately defined [^1] We will define this phase here relative to the (non-perturbative) chiral coupling $G_{8}$. In order to correctly compute the interference term in the $K_{L}\to{\pi^{0}}\ell^{+}\ell^{-}$ branching fraction knowledge of the relative phase between $G_{8}$ and the imaginary part of the perturbative parameter $C_{7V}$ is also necessary.) phase of the $K_{S}\to{\pi^{0}}\ell^{+}\ell^{-}$ amplitude from experiment, given a theoretical framework for analysing the data. The framework that has been usually employed is the so-called “beyond one-loop model” (B1L) proposed in ref. [17]. As compared with the calculation in the chiral expansion at order $p^{4}$ [18, 19] the B1L model includes terms of higher chiral order, which are needed in order to properly describe the experimental energy dependence. The B1L amplitudes are written as follows
@@ 3.1 Asymptotic behaviour of $W^{[1/2]}(s;\nu)$ and $W^{[3/2]}(s;\nu)$
$\lim_{q\to\infty}\int d^{4}x\,e^{iqx}T\{j_{\mu}^{em}(x)\mathcal{L}_{\Delta{S}=1}(0)\}=(q_{\mu}q_{\nu}-g_{\mu\nu}q^{2})(A+B\ln\frac{-q^{2}}{\nu^{2}}+\cdots)\,[\bar{s}\gamma_{\nu}(1-\gamma^{5}){d}]$ (17)
@@
$\begin{array}[]{ll}{\displaystyle\int}d^{4}x\,e^{iq\cdot x}T\{j^{\mu}(x){\cal L}^{eff}_{\Delta{S=1}}(0)\}_{\rm dim=6}\sim&\bigg[C^{(6)}_{V}(\nu)\dfrac{g^{\mu\beta}q^{\alpha}-g^{\mu\alpha}q^{\beta}}{q^{2}}\\ &+C^{(6)}_{A}(\nu)\epsilon^{\mu\nu\alpha\beta}\dfrac{q_{\nu}}{q^{2}}\bigg]\times(\bar{s}\gamma_{\alpha}{u})(\bar{u}\gamma_{\beta}{d})\ {.}\end{array}$ (18)
@@ 3.2 Muskhelishvili-Omnès representations
$\Omega^{I}_{1}(s)=\exp\left[\frac{s}{\pi}\int_{m_{+}^{2}}^{\infty}ds^{\prime}\dfrac{\delta_{1}^{I}(s^{\prime})}{s^{\prime}(s^{\prime}-s)}\right]\$(19)
where $m_{+}=m_{\pi}+m_{K}$. The $K\pi$ phases $\delta_{1}^{I}(s)$ to be used in these integrals must be equal to the elastic scattering phase-shifts in the elastic region and can be chosen in a somewhat arbitrary way above that region. It is noteworthy that the elastic region extends effectively well above the first inelastic threshold $K\pi\pi$, up to $\Lambda=1.3$ GeV approximately [37]. The $J=1$, $I=3/2$ phase-shift is very small and will be simply interpolated to zero above 1 GeV. In the case of $I=1/2$ we can use the $K\pi$ vector form factor (the modulus of which is a measurable quantity) i.e. take $\Omega_{1}^{1/2}(s)\equiv f_{+}^{K\pi}(s)/f_{+}^{K\pi}(0)$ which obeys eq. (19) with a phase given by the $K\pi$ form factor phase in the inelastic region.
@@
$W^{[1/2]}(s;\nu)=W^{[1/2]}_{\pi\pi}(s)+f_{+}^{K\pi}(s)\Big[C^{[1/2]}(\nu)-\frac{s}{\pi}\int_{m_{+}^{2}}^{\Lambda^{2}}ds^{\prime}\frac{W^{[1/2]}_{\pi\pi}(s^{\prime}){\rm Im\,}[1/f_{+}^{K\pi}(s^{\prime})]}{s^{\prime}(s^{\prime}-s)}\Big]\ .$ (24)
@@
$W^{[1/2]}(s)=W^{[1/2]}_{\pi\pi}(s)+f_{+}^{K\pi}(s)\Big[G_{F}m_{K}^{2}\frac{2a_{+}-a_{S}}{f_{+}^{K\pi}(0)}-\frac{s}{\pi}\int_{m_{+}^{2}}^{\Lambda^{2}}ds^{\prime}\frac{W^{[1/2]}_{\pi\pi}(s^{\prime}){\rm Im\,}[1/f_{+}^{K\pi}(s^{\prime})]}{s^{\prime}(s^{\prime}-s)}\Big]\$(25)
@@ 4.2 Input for the $K\pi\to\pi\pi$ amplitudes
$\widetilde{M}_{1}(s)=\tilde{\mu}_{1}\widetilde{{\cal S}}_{1}(s)\ ,\quad\widehat{\widetilde{M}}_{1}(s)=\frac{\tilde{\mu}_{1}}{2\kappa(s)}\int_{-1}^{+1}dz(-9(s-s_{0})z-3\kappa(s)z^{2})\widetilde{S}_{1}(t(z))\$(35)
@@ 5.5 Additional contributions: $\omega$ and $\phi$ resonances
$\bar{b}_{+}|_{\omega,\phi}\simeq+0.25\ (a_{S}>0),\quad\bar{b}_{+}|_{\omega,\phi}\simeq+0.26\ (a_{S}<0)\$(54)
using $M_{V}=M_{\omega}$. This reflects the fact that ${\rm Im\,}[W_{+}]$ (unlike ${\rm Re\,}[W_{+}]$), from which the combination $2g_{V}^{5}-g_{V}^{6}$ is derived, has a very small dependence on $a_{S}$ at low energy and up to the $\rho$ mass.
@@
The results of including the effects of the $I=0$ resonances, estimated in the way described above, is illustrated in fig. 13. In the case of $W_{+}$, they decrease the modulus squared of the amplitudes in both cases of $a_{S}>0$ and $a_{S}<0$. The former case thus remains favoured by comparison with the experimental data. A similar effect of the $I=0$ resonances to $W_{+}$ was observed also in the factorisation approach [29, 65]. In the case of $|W_{S}|^{2}$, the $I=0$ resonances modify significantly the difference between the results corresponding to $a_{S}>0$ and $a_{S}<0$ (see fig. 13, right).
@@ 5.6 Additional contributions: higher mass resonances
$C^{[1/2]}(s;\nu)=\frac{G_{F}}{\sqrt{2}}V_{ud}V_{us}\frac{8m_{K}^{2}}{3}\,(N_{c}(C_{1}-C_{4})+C_{2}-C_{3})\Big(\frac{5}{3}+\log\frac{\nu^{2}}{M^{2}}-\psi(3-\frac{s}{M^{2}})\,\Big)\ .$ (55)
@@ 6 Conclusions
Other light vector resonances are present, in particular $\omega(782),\phi(1019)$, but their influence on the $W_{+,S}$ amplitudes cannot be accounted for in the same way as was done for the $\rho$ and the $K^{*}$ resonances, because the corresponding discontinuities involve amplitudes, $\pi\to 3\pi, KK$${\bar{K}}\to{K}\pi$ which cannot be accessed experimentally. We estimated their contributions in an approximate way by appealing to nonet symmetry, $\Delta{I}=1/2$ dominance, and relying on a chiral-resonance Lagrangian, finding them to be less important than the $I=1,1/2$ resonances but possibly significant. In a general way, the isoscalar resonances can be accounted for model independently by introducing one additional subtraction parameter in the $W^{[1/2]}$ dispersive representation.